Mang Berto, when asked about his age, replies, “I am five years older than twice the age of my youngest child” .Express his age in algebraic equation if his age is 55.
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

let x = the age of his youngest child.
then 55 = 2 * x + 5.
that's your equation.
solve for x to get:
x = (55 - 5) / 2 = 50 / 2 = 25.
his son's age is 25.
he is twice that plus 5 = 55.
the general equation is y = 2 * x + 5.
if you know his son's age, then you solve for y.
if you know his age, then you solve for x.

Answer by ikleyn(52781)

.

The equation is as you read the problem

    55 = 2x + 5,


where x is the age of the youngest child.

Solved / answered.